OpenStudy (anonymous):

Rewrite the equation to isolate the variable y=9x+2 and df=g-10

OpenStudy (bahrom7893):

isolate which variable?

OpenStudy (anonymous):

oh im sorry , for the first one its for x and the second one for d

OpenStudy (bahrom7893):

1) subtract 2 from both sides: y-2 = 9x +2 -2 y-2 = 9x 2) divide both sides by 9: (y-2)/9 = 9x/9 (y-2)/9=x

OpenStudy (bahrom7893):

Do the 2nd one yourself. I'll help you if you're stuck.

OpenStudy (anonymous):

okay..im coming let me try to solve it.

OpenStudy (anonymous):

so first I subtracted 10 from both sides... df=g-10 +10 +10 the 10s cancel out... then that leaves me with df=g ... I think ? am I doing it right so far?

OpenStudy (bahrom7893):

wait wrong

OpenStudy (anonymous):

no , I add 10 to both sides..

OpenStudy (bahrom7893):

you want to isolate d

OpenStudy (bahrom7893):

so you just need to perform the operations that will clean up the side that d is on

OpenStudy (anonymous):

so , don't add 10

OpenStudy (anonymous):

?

OpenStudy (bahrom7893):

no, divide both sides by f

OpenStudy (anonymous):

okay so then , d=g-10/f ?

OpenStudy (bahrom7893):

well yea but it would be: (g-10)/f

OpenStudy (anonymous):

okay, thank you. may I get your help on another one like these?

OpenStudy (bahrom7893):

sure

OpenStudy (anonymous):

1/3s-8=t solve for s

OpenStudy (bahrom7893):

can you use parenthesis for that please? im confused about what the equation is.

OpenStudy (anonymous):

there are none

OpenStudy (bahrom7893):

i mean can you format the question properly?

OpenStudy (anonymous):

 is that better ?